B , C and D vying for for the position.



List all the possible combinations of applicants to be hired.



Use the formula for combination to solve for the number of possible of



filling the vacant positions.




1
Expert's answer
2021-12-01T18:09:50-0500

there are 3 vacant positions and 4 applicants A, B, C, D vying for for the position


all the possible combinations of applicants to be hired:

"(A,B,C),(A,B,D),(A,C,D),(B,C,D)"


number of possible of filling the vacant positions is number to take 3 items from total 4:

"n=C^3_4=\frac{4!}{3!1!}=4"